Calculate the amount and compound interest on 
₹ 18000 for 2 whole 𝟏/ 𝟐 years at 10% per annum compounded annually

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Principal (P) = ₹ 18,000
Rate (R) = 10% annual
Number of years (n) = 2 whole 1/2
The amount for 2 years and 6 months can be calculated by calculating the amount for 2 years using the compound interest formula, and then calculating the simple interest for 6 months on the amount obtained at the end of 2 years.
First, the amount for 2 years has to be calculated.
Amount, A =P (1 +R/100 )^n
=₹[18000(1+1/10)^2]
=₹[18000(11/10)^2]
=₹21780
By taking ₹ 21780 as principal, the S.I. for the next1/2year will be calculated.
S.I =21780×1/2×10=₹1089
100
∴ Interest for the first 2 years = ₹ (21780 − 18000) = ₹ 3780
And interest for the next1/2year = ₹ 1089
∴Total C.I.= ₹ 3780 + ₹ 1089 = ₹ 4,869
Amount, A = P + C.I.
= ₹ 18000 + ₹ 4869
= ₹ 22,869

Thanks